### v1

# Ground-based Image Deconvolution with Swin Transformer UNet

Abstract
Aims. As ground-based all-sky astronomical surveys will gather millions of images in the coming years, a critical requirement emerges for the development of fast deconvolution algorithms capable of efficiently improving the spatial resolution of these images. By successfully recovering clean and high-resolution images from these surveys, our objective is to help deepen our understanding of galaxy formation and evolution through accurate photometric measurements.
Methods. We introduce a two-step deconvolution framework using a Swin Transformer architecture. Our study reveals that the deep learning-based solution introduces a bias, constraining the scope of scientific analysis. To address this limitation, we propose a novel third step relying on the active coefficients in the sparsity wavelet framework.
Results. By conducting a performance comparison between our deep learning-based method and Firedec, a classical deconvolution algorithm, we analyze a subset of the EDisCS cluster samples. We demonstrate the advantage of our method in terms of resolution recovery, generalization to different noise properties, and computational efficiency. Not only does the analysis of this cluster sample assess the efficiency of our method, but it also enables us to quantify the number of clumps within these galaxies in relation to their disc colour. This robust technique holds promise for identifying structures in the distant universe from ground-based images.
Key Words.:
Deconvolution – Denoising – Swin Transformer – SUNet – VLT – HST – Clump Detection1 Introduction
High spatial resolution and high signal-to-noise observations are prerequisites to most observational astrophysical problems. However, expecting the two conditions to happen simultaneously is challenging: space telescopes have limited collecting power, while large telescopes are ground-based and are therefore affected by atmospheric turbulence. This is clearly illustrated by the two main missions/observatories of the next decade: the ESA-NASA Euclid space telescope (Euclid Collaboration et al., 2022; Laureijs et al., 2011) and the Vera C. Rubin Observatory (Ivezić et al., 2019). Exploiting the best of both worlds is possible, provided reliable post-processing techniques are developed to remove blurring by the atmosphere and instrument Point Spread Function (PSF). To further complicate matters, sensor variations introduce noise in the images. Hence, image deconvolution in astrophysics is an ill-posed and ill-conditioned inverse problem that requires regularization to achieve a unique solution. This was realized very early in the field, and solutions were proposed, such as minimizing the Tikhonov function (e.g. Tikhonov & Arsenin, 1977) or maximizing the entropy of the solution (e.g. Skilling & Bryan, 1984). Other algorithms, based on Bayesian statistics, include the Lucy-Richardson algorithm, used on the early data from the Hubble Space Telescope (HST) (Richardson, 1972; Lucy, 1974). Magain et al. (MCS; 1998) proposed a two-channel method that separates the point sources from the spatially extended ones and that deconvolves using a narrow PSF, hence achieving finite but improved resolution well compatible with the sampling (pixel) chosen to represent the solution. Improvements of the ”MCS” method implement wavelet regularization of the extended channel of the solution Cantale et al. (2016a). This method was further refined by Michalewicz et al. (STARRED; 2023) who use an isotropic wavelet basis, called Starlets (Starck et al., 2015), to regularize the solution.
However, deep learning offers a completely different approach, by learning the properties of the desired solution. Once trained, deep learning-based methods are also orders of magnitude faster than classical methods. Of note, UNets (Ronneberger et al., 2015) have become popular due to their highly non-linear processing and multi-scale approach. Building on Unet, Sureau et al. (2020) developed the Tikhonet method for deconvolving galaxy images in the optical domain. They demonstrated that Tikhonet outperformed sparse regularization-based methods in terms of Mean Squared Error (MSE) and a shape criterion, where a measure of the galaxy ellipticity was used to encode its shape. Nammour et al. (2022) added a shape constraint to the Tikhonet loss function and achieved better performance. In our recent work (Akhaury et al., 2022), we proposed a new deconvolution approach that employs the Learnlet decomposition (Ramzi et al., 2023), using the same two-step approach as in Sureau et al. (2020), but by substituting the Unet denoiser by Learnlet.
With the recent advent of Vision Transformers (Dosovitskiy et al., 2021), significant progress has been made in the field of image restoration (Liang et al., 2021; Zamir et al., 2022; Wang et al., 2022). This is the motivation for the present work to investigate the performance of SUNet (Fan et al., 2022), a variant of Unet with Swin Transformer blocks (Liu et al., 2021) replacing the convolutional layers. To our knowledge, SUNet has not yet been used as a denoiser in a deconvolution framework. We show that the neural network outputs introduce bias, thereby limiting the scientific analysis. The bias appears as a small flux loss in the outputs. To counter this, we propose a third debiasing procedure based on the active coefficients in the sparsity wavelet framework, called Multi-resolution support (Starck et al., 1995), as explained in Section 2.3. Our experiments involved real HST images, with the network trained on images extracted from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011).
Finally, to assess the generalizability, we test our new code on a completely different sample of ground-based images obtained with the FORS2 camera at the Very Large Telescope (VLT). Spatially resolved images of galaxies can help address a multitude of topics involving their morphologies. In the present case, we tackle the question of how the galaxy cluster environment could impact the properties of the discs of spiral galaxies. From the same dataset, Cantale et al. (2016b) had found that of cluster spiral galaxies at redshift - had disc colours redder by more than of the mean colours of their field counterparts. This result was obtained thanks to the deconvolution code Firedec (Cantale et al., 2016a). The VLT and -band images, with initial spatial resolutions between and had been deconvolved with a target final resolution of on pixels. Even though the gain in resolution was indeed substantial, conditions were not sufficient to go beyond the global photometric properties of the discs. In this study, we go one step further and investigate their internal structure, specifically by identifying their star-forming clumps.
Compact star-forming clumps are identified in distant galaxies, particularly when they benefit from HST deep images (e.g., Wuyts et al., 2012; Guo et al., 2015; Sattari et al., 2023). They are understood as playing a crucial role in galaxy assembly and star-formation activity. Recent research by Sok et al. (2022) explored clump fractions in star-forming galaxies (SFGs) from multi-band analysis in the COSMOS field. The match between HST and ground-based resolution was performed with Firedec. Their findings indicated a decline in the fraction of clumpy galaxies with increasing stellar masses and redshifts. Moreover, they observed that clumps contributed a higher fractional mass towards galaxies at higher redshifts. In our study, by employing a more powerful deconvolution algorithm capable of accurately recovering small-scale structures at high spatial resolution from ground-based multi-band observations (as demonstrated in Section 4.2), our goal is to quantify the number of clumps in EDisCS cluster galaxies and examine their relationship with disc colour.
In Section 2, we present the deconvolution problem and introduce our proposed deep learning method to address it. The process of generating our dataset and conducting experiments is outlined in Section 3. In Section 4, we present the results of our deconvolution algorithm. Finally, we draw our conclusions in Section 5. To support reproducible research, the codes and trained models utilized in this article are publicly accessible in Section 6. Additional studies and supplementary information can be found in Appendices A, B, and C.
2 Deep-Learning based Deconvolution
The deconvolution problem can be summarized in a very simple (but hard to solve) equation. Let be the observed image and be the PSF. The observed image can be modelled as
where denotes the target image, denotes the convolution operation, and denotes additive Gaussian noise. The goal is then to recover the ground truth image , given the PSF convolution and the unknown noise. Such ill-posed inverse problems require regularization of the solution, to select the most appropriate one among the many others that are compatible with the data. Sparse wavelet regularization using the or norm has remained the most accepted one in the past, but the recent advent of machine learning methods has changed this paradigm.
2.1 Tikhonov Deconvolution
Tikhonov Deconvolution is a two-step deep learning-based deconvolution technique. In the first step, the input images undergo deconvolution using a Tikhonov filter with quadratic regularization. If denotes the circulant matrix associated with the convolution operator , the Tikhonov solution of equation 1 is expressed as:
Here, represents the linear Tikhonov filter, configured as a Laplacian high-pass filter to penalize high frequencies. The regularization weight, denoted as , is determined through a grid search. The Tikhonov step produces deconvolved images containing correlated additive noise, which is subsequently eliminated in the following step by an appropriate denoiser. The denoisers are trained to learn the mapping from the Tikhonov output to the ground truth image by minimizing a suitable loss function, such as or .
The denoising performance is significantly influenced by the choice of the model architecture. To effectively capture distant correlations, it is crucial to incorporate multi-scale processing in the model design. This consideration leads to the adoption of a layout similar to that of a UNet (Ronneberger et al., 2015). Additionally, Mohan et al. (2020) demonstrated that biases in convolutional layers can result in low generalization capability. Consequently, for our experiments, we opt for bias-free networks.
2.2 SUNet Denoising
In recent times, the UNet architecture has become popular in various image-processing applications due to its incorporation of hierarchical feature maps, facilitating the acquisition of multi-scale contextual features. It is widely employed in diverse computer vision tasks, including segmentation and restoration (Yu et al., 2019; Gurrola-Ramos et al., 2021). Evolved versions such as Dense-UNet (Guan et al., 2020), Res-UNet (nan Xiao et al., 2018), Non-local UNet (Yan et al., 2020), and Attention UNet (Jin et al., 2020) have also emerged. Thanks to its flexible structure, the UNet can adapt to different building blocks, enhancing its overall performance. Moreover, the evolution of image processing methodologies has seen the introduction of Transformer models (Vaswani et al., 2017), which were initially successful in natural language processing (NLP) but have also demonstrated impressive performance in image classification (Dosovitskiy et al., 2021; Yuan et al., 2021). However, when directly applied to vision tasks, transformers face challenges such as the large-scale difference between images and sequences, making them less effective in modelling long sequences due to the need for a square number of parameters for 1-dimensional sequences. Additionally, transformers are not well-suited for dense prediction tasks such as instance segmentation at the pixel level. Swin Transformer addresses these issues by introducing a shifted-window mechanism to reduce the number of parameters, establishing itself as a state-of-the-art solution for high-level vision tasks (Liu et al., 2021). Taking inspiration from these advancements, Fan et al. (2022) incorporated Swin Transformers as building blocks within the UNet architecture and showed that they could achieve competitive results compared to existing benchmarks for image denoising. The network architecture is visually depicted in Figure 1, and the PyTorch code is publicly available on GitHub (details in Section 6).
While SUNet was originally developed for White Gaussian Noise (WGN) removal, the Tikhonov deconvolution step (equation 2) introduces alterations to the noise characteristics in the images. Consequently, it becomes crucial to assess the extent to which SUNet can effectively handle the presence of Colored Gaussian Noise (CGN). Thus, our tests examine SUNet’s generalizability in the presence of CGN.
2.3 Debiasing with Multi-resolution Support
Our study reveals that the deep learning-based solution introduces a bias, evident in the form of positive structures in the residuals, as illustrated in Figure 2(a). This bias can have implications for the accuracy of scientific analyses, potentially influencing the flux estimation within the recovered features in the reconstructed images. To recover the lost flux and enhance image sharpness, multi-resolution support has proven effective (Starck et al., 1995). Denoting the Starlet transform as and the SUNet output solution as , an iterative process allows the recovery of flux from the residual . In each iteration, denoted as , a debiasing correction term is applied by multiplying the multi-resolution support matrices of the SUNet output solution at each scale with the Starlet decomposition of the gradient of the residual, incrementally modifying the deconvolved image:
where
The acronym MRS stands for multi-resolution support, indicating a boolean measure of whether an image contains information at a specific pixel and scale. If represents the wavelet coefficient at a given scale and is the threshold value, the multi-resolution support operation can be defined as:
The flux-recovery process is illustrated in Figure 2. The iterative process is stopped once convergence is achieved in the standard deviation of the residual as a function of the number of iterations, as seen in sub-figure 2(d). A more detailed study on the impact of debiasing with multi-resolution support on neural networks is presented in Appendix B.
3 Dataset and Experiments
3.1 Training Dataset Generation
We extracted Hubble Space Telescope (HST) cutouts measuring pixels from CANDELS (Grogin et al., 2011; Koekemoer et al., 2011) in the filter (-band). These cutouts were then convolved with a Gaussian PSF having a Full Width at Half Maximum (FWHM) of 15 pixels. Following the convolution, we injected White Gaussian Noise (WGN) with a standard deviation denoted as . This choice ensured that the faintest object in our dataset had a peak Signal-to-Noise Ratio (SNR) close to and was barely visible. With this particular value of , our dataset exhibited a range of SNR values depending on the magnitude of each galaxy. To standardize the images, each one was normalized within the range. This normalization involved subtracting its own mean and scaling its peak value by the maximum value in the image. Finally, the batch of images was randomly divided into Training-Validation-Test subsets in the ratio .
3.2 Training the SUNet
The SUNet architecture was trained using a Titan RTX Turing GPU with 24 GB RAM for each job. The training aimed at learning the mapping from the Tikhonov output (containing CGN) to the corresponding HST image . The training utilized the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of , gradually halving every 25 epochs until reaching a minimum of . As in the original SUNet paper (Fan et al., 2022), the -loss was used for training. A more detailed discussion on the significance of the training loss function is given in Appendix C. The input images were processed in mini-batches of size . The dataset was augmented with random rotations in multiples of 90°, translations and flips along horizontal & vertical axes. Starting with images in our initial training dataset, we increase its diversity by a factor of by performing augmentation.
3.3 Test Dataset
We utilized the ESO Distant Cluster Survey, EDisCS (White, S. D. M. et al., 2005), as our benchmark dataset to assess the performance of our deconvolution method. EDisCS is an extensive ESO Large Programme focused on the analysis of galaxy clusters within the redshift range , covering a diverse range of masses with velocity dispersions ranging from approximately km s-1 to km s-1. All clusters benefit from deep , , , and photometry obtained with FORS2 at the VLT. Additionally, a subset of ten clusters was imaged with the ACS at the HST in the F814W filter (Simard et al., 2002). Previous work by Cantale et al. (2016b) employed the deconvolution technique Firedec to analyze spiral disc colours for EDisCS cluster galaxies, investigating trends with cluster masses and lookback-time. In our study, we focus on analyzing a subset of EDisCS clusters at three distinct redshifts—, , and —using our proposed deconvolution method, and go a step further by investigating the star-forming clumps in these galaxies. Table 1 provides a summary of the properties of these clusters and the number of galaxies in each of them.
| Cluster ID | ||
| cl1037.9-1243 | 0.5805 | 11 |
| cl1040.7-1155 | 0.7043 | 19 |
| cl1054.4-1146 | 0.6972 | 20 |
| cl1103.7-1245b | 0.7031 | 6 |
| cl1216.8-1201 | 0.7955 | 28 |
Our analysis specifically considered the , , and photometric bands, allowing us to evaluate the effectiveness of our deconvolution method in capturing variations across these photometric bands. To enable a more detailed analysis, we grouped these EDisCS galaxies based on their disc colour in order to study the different trends. Since the EDisCS clusters were solely observed in the F814W filter for HST, our deconvolution method presents a unique opportunity to extract high spatial resolution galaxy properties from the ground-based FORS2 multi-band observations. Furthermore, we demonstrate the superiority of SUNet over Firedec in generating cleaner deconvolved images, as detailed in Sections 4.2 and 4.3.
4 Results
The EDisCS images in the , , and bands served as inputs for the SUNet deconvolution framework, yielding corresponding deconvolved outputs. Remarkably, the algorithm operates with exceptional speed, requiring only ms to deconvolve a single image on a Titan RTX Turing GPU with 24 GB RAM—approximately times faster than Firedec on average. A histogram of the magnitudes of the same objects in the HST F814W filter is shown in Figure 3. In Section 4.1, we detail our approach to detecting sizes and the number of clumps in the deconvolved outputs. Section 4.2 then presents a comparative analysis of the quality of our SUNet outputs against Firedec (Cantale et al., 2016a). With the validity of our method established, Section 4.3 presents the deconvolved results, followed by a thorough analysis in Section 4.4.
4.1 Objects Size and Clump Detection
To detect clumps, or small-scale structures, and measure the sizes of galaxies, we employed the SCARLET Python package (Melchior et al., 2018). Utilizing the Starlet transform from SCARLET, our approach involved decomposing images into different scales, where each scale captures a specific frequency component. Ensuring consistency for an unbiased comparison, we maintained fixed algorithm parameters across different bands and objects. All images underwent decomposition into five scales, with the fourth scale chosen for size detection and the second scale for clump detection. A detection threshold was applied to each scale during the process. In the context of clump detection, the Starlet transform was computed using the standard deviation solely within the region enclosed by the size detection outline, rather than considering the entire image. This refined approach ensured more precise thresholding in the Starlet space. Finally, a clump was only considered valid if it lay within the size detection outline, ensuring that background artefacts were excluded. In Figure 4, we present examples of size detection and their corresponding clump detection cases.
4.2 Comparison with Classical Methods
We conducted a thorough performance comparison between SUNet and Firedec, a classical deconvolution method based on wavelet regularization (Cantale et al., 2016a). For direct comparison with HST quality, we concentrate on the -band outputs for each method, as the EDisCS clusters were exclusively observed in the F814W filter for HST. Both methods exhibit better performance on low-magnitude (or high-SNR) images, with a gradual decline in performance in the high-magnitude (or low-SNR) regime. In this case, the Mean Squared Error (MSE) metric between the deconvolved outputs and the ground truth HST images is not a robust metric for indicating similarity, since it would be biased by the background noise in the HST images (Wang & Bovik, 2009). Instead, we use the Structural Similarity Index (SSIM), a full reference metric which quantifies the similarity between two images by comparing their structural information or spatial inter-dependencies (Wang et al., 2004). An SSIM of 1 implies identical images. The observed trends are illustrated in Figure 5. We further assessed the ability of the deconvolution algorithms to accurately resolve small-scale structures. For this, we leveraged the SCARLET Python package, as detailed in Section 4.1. The fraction of area overlap between small-scale structure detections in the deconvolved outputs and HST is depicted in Figure 6. Based on these metrics, SUNet clearly outperforms Firedec.
Building on the foundation of Firedec, an enhanced method named STARRED was recently introduced by Michalewicz et al. (2023). STARRED brings innovation by incorporating an isotropic wavelet basis known as Starlets (Starck et al., 2015) to refine the regularization process in solving the deconvolution problem. The outputs of Firedec, STARRED, and SUNet are shown in Figure 7. Upon visual inspection of the outputs and the residuals (residual noisy image PSF deconvolved image), it is evident that SUNet consistently generalizes better than Firedec and STARRED.
4.3 SUNet Deconvolution Results
The methods were put to the test using real ground-based images captured by the FORS2 camera at the Very Large Telescope (VLT) in Chile. Notably, SUNet showed a remarkable ability to generalize effectively to images with entirely different noise properties than those present in the training dataset, as depicted in Figure 8. As illustrated, we are successfully able to recover the morphology and lost small-scale structures. Figure 9 presents the trend in the relative flux error between the SUNet deconvolved outputs and the corresponding HST targets as a function of the total clump size. The uncertainty in the flux level is higher for smaller clumps.
4.3.1 Resolution Recovery
To gauge the achieved resolution in the deconvolved outputs, we calculated the average ratio between the areas of the smallest detected clump in the SUNet output and its counterpart in the HST image. This ratio, approximately , implies an average SUNet output resolution of around , considering the known HST resolution of .
4.3.2 False Positives and False Negatives
To assess the reliability of our deconvolution method for real-world applications, we conducted an analysis to estimate the number of false positives and false negatives in our study. We ran SCARLET to detect clumps in both SUNet outputs and HST ground truths. For each HST clump, we checked whether the SUNet-identified clump centroid fell within a -pixel radius of the HST clump centroid. Clumps failing this criterion were considered false positives. The false positive rate was computed across our entire EDisCS dataset by tallying the total count of false positives and dividing it by the overall count of detected clumps in the HST images. However, it is important to note that this result may be biased due to SCARLET’s performance. Instances exist where SCARLET identifies a clump in the SUNet image but misses it in the corresponding HST image, leading to an elevated false positive count. To address this, we employed visual inspection to filter out falsely detected cases. The resultant false positive rate was determined to be approximately . Using a similar approach, we also computed the false negative rate, indicating the probability of missing clumps in the deconvolved images that are present in the HST image. This rate was found to be , signifying a very low probability of missing features.
4.4 Analysis of Deconvolved EDisCS Galaxies
As a sanity check, we verified that any conclusions drawn regarding the internal properties of our sample galaxies are not influenced by biases in their sizes, either in relation to redshift or disc colours. To this end, the histogram of galaxy sizes grouped by their parent cluster redshift and disc colour is presented in Appendix A. Both plots show that all galaxies have the same global spatial extent.
Following Cantale et al. (2016b), our analysis focuses on the population of galaxies with discs redder than their field counterparts. Some of our sample galaxies have normal colours, hence falling in the colour distribution of the field galaxies, some are bluer (by more than of the colour distribution). A few known physical processes can induce enhanced star-forming activity, but we are instead interested in the possible evidence for quenching mechanisms. Therefore, in the latter case, normal and blue disc galaxies form a common broad class of systems to which we compare the redder ones. Employing the clump detection method outlined in Section 4.1, we computed the histogram of the number of clumps in galaxies, categorized by disc colour, as shown in Figure 10. Note that the only-1-clump case reflects the identification of the central and luminous part (bulge) of the galaxies. As illustrated in Figures 4 and 8, clumps in the -band are brighter than those in the and -bands, in agreement with the spectral energy distribution of young stellar populations. In principle, it is therefore easier to detect clumps in the -band at equivalent photometric depths of the images. This may explain the more continuous distribution in the number of clumps, from to , in the -band. Even so, as witnessed in Figure 10, the general trend is the same from one band to the other. This trend is clear and reveals that the red discs, that were initially identified by Cantale et al. (2016b), have fewer clumps than their bluer counterparts, most likely due to an earlier cessation of star formation. This result opens promising prospects for future studies on larger samples and over larger look-intervals.
5 Conclusion
We proposed a deconvolution framework involving a two-step process: Tikhonov deconvolution, post-processing with an SUNet denoiser, and an additional debiasing step using multi-resolution support. SUNet was trained on galaxy images from the CANDELS survey and demonstrated superior performance compared to Firedec in the astrophysical context. After establishing the validity of our method, we applied it to deconvolve a set of galaxies from the EDisCS cluster at three different redshifts. Using SCARLET, we provided further analysis of the galaxies in terms of their size and disc colour. We quantified the number of clumps in these galaxies, examining their relationship with disc colour. Our results, based on both quantitative metrics and visual assessments, highlighted the effectiveness of SUNet, showcasing its ability to generalize well to unseen real images with diverse noise properties, which can be attributed to its transformer-based backbone involving the self-attention mechanism.
In summary, this work introduces and evaluates an advanced deconvolution framework applied to ground-based astronomical images. The key findings and contributions include:
Resolution recovery: Based on our SCARLET detection procedure, SUNet demonstrates the capability to recover small-scale structures with an average resolution of approximately and outperforms classical algorithms like STARRED and Firedec (Section 4.3.1).
Generalization to diverse noise properties: The method showcases robust generalization to noise properties different from its training dataset, indicating its adaptability to various observational conditions.
Clump analysis: Red discs exhibit fewer clumps than their bluer counterparts, affirming the lower presence of star-forming regions.
False positive and False negative rates: Based on our SCARLET detection analysis on EDisCS, SUNet maintains a false positive rate of and a false negative rate of , ensuring reliable feature recovery (Section 4.3.2).
Computational efficiency: The AI-based framework proves to be highly efficient, with an execution time of approximately ms per image, making it around times faster than traditional deconvolution methods like STARRED and Firedec.
Hence, this technique can be used to efficiently identify structures in the distant universe at high spatial resolution from ground-based images. The technique’s applicability to multi-band observations further enhances its utility in studying various astrophysical phenomena. The efficiency of SUNet in processing large datasets and accelerating the deconvolution process opens up opportunities for swift analyses. Access to such a fast and robust deconvolution framework holds the potential to facilitate numerous astrophysical investigations.
6 Reproducible Research
[…]
1.
[…]
2.
[…]
3.
[…]
4.
[…]
Acknowledgements.
This work was funded by the Swiss National Science Foundation (SNSF) under the Sinergia grant number CRSII5_198674. This work was supported by the TITAN ERA Chair project (contract no. 101086741) within the Horizon Europe Framework Program of the European Commission, and the Agence Nationale de la Recherche (ANR-22-CE31-0014-01 TOSCA). The authors thank David Donoho for useful discussions.References
- Akhaury et al. (2022) Akhaury, U., Starck, J.-L., Jablonka, P., Courbin, F., & Michalewicz, K. 2022, Frontiers in Astronomy and Space Sciences, 9
- Cantale et al. (2016a) Cantale, N., Courbin, F., Tewes, M., Jablonka, P., & Meylan, G. 2016a, A&A, 589, A81
- Cantale et al. (2016b) Cantale, N., Jablonka, P., Courbin, F., et al. 2016b, A&A, 589, A82
- Dosovitskiy et al. (2021) Dosovitskiy, A., Beyer, L., Kolesnikov, A., et al. 2021, in International Conference on Learning Representations
- Euclid Collaboration et al. (2022) Euclid Collaboration, Scaramella, R., Amiaux, J., et al. 2022, A&A, 662, A112
- Fan et al. (2022) Fan, C.-M., Liu, T.-J., & Liu, K.-H. 2022, in 2022 IEEE International Symposium on Circuits and Systems (ISCAS) (IEEE)
- Grogin et al. (2011) Grogin, N. A., Kocevski, D. D., & Faber, S. M. 2011, The Astrophysical Journal Supplement Series, 197, 35
- Guan et al. (2020) Guan, S., Khan, A. A., Sikdar, S., & Chitnis, P. V. 2020, IEEE Journal of Biomedical and Health Informatics, 24, 568
- Guo et al. (2015) Guo, Y., Ferguson, H. C., Bell, E. F., et al. 2015, ApJ, 800, 39
- Gurrola-Ramos et al. (2021) Gurrola-Ramos, J., Dalmau, O., & Alarcón, T. E. 2021, IEEE Access, 9, 31742
- Ivezić et al. (2019) Ivezić, Ž., Kahn, S. M., Tyson, J. A., et al. 2019, ApJ, 873, 111
- Jin et al. (2020) Jin, Q., Meng, Z., Sun, C., Cui, H., & Su, R. 2020, Frontiers in Bioengineering and Biotechnology, 8
- Kingma & Ba (2014) Kingma, D. P. & Ba, J. 2014 [arXiv:1412.6980]
- Koekemoer et al. (2011) Koekemoer, A. M., Faber, S. M., & Ferguson, H. C. 2011, The Astrophysical Journal Supplement Series, 197, 36
- Laureijs et al. (2011) Laureijs, R., Amiaux, J., Arduini, S., et al. 2011, arXiv e-prints, arXiv:1110.3193
- Liang et al. (2021) Liang, J., Cao, J., Sun, G., et al. 2021, in Proceedings of the IEEE/CVF international conference on computer vision, 1833–1844
- Liu et al. (2021) Liu, Z., Lin, Y., Cao, Y., et al. 2021, CoRR, abs/2103.14030 [2103.14030]
- Lucy (1974) Lucy, L. B. 1974, AJ, 79, 745
- Magain et al. (1998) Magain, P., Courbin, F., & Sohy, S. 1998, The Astrophysical Journal, 494, 472
- Melchior et al. (2018) Melchior, P., Moolekamp, F., Jerdee, M., et al. 2018, Astronomy and Computing, 24, 129
- Michalewicz et al. (2023) Michalewicz, K., Millon, M., Dux, F., & Courbin, F. 2023, Journal of Open Source Software, 8, 5340
- Mohan et al. (2020) Mohan, S., Kadkhodaie, Z., Simoncelli, E. P., & Fernandez-Granda, C. 2020, in International Conference on Learning Representations
- Nammour et al. (2022) Nammour, F., Akhaury, U., Girard, J. N., et al. 2022, A&A, 663, A69
- nan Xiao et al. (2018) nan Xiao, X., Lian, S., Luo, Z., & Li, S. 2018, 2018 9th International Conference on Information Technology in Medicine and Education (ITME), 327
- Ramzi et al. (2023) Ramzi, Z., Michalewicz, K., Starck, J.-L., Moreau, T., & Ciuciu, P. 2023, Journal of Mathematical Imaging and Vision, 65, 240
- Richardson (1972) Richardson, W. H. 1972, Journal of the Optical Society of America (1917-1983), 62, 55
- Ronneberger et al. (2015) Ronneberger, O., Fischer, P., & Brox, T. 2015, CoRR, abs/1505.04597 [arXiv:1505.04597]
- Sattari et al. (2023) Sattari, Z., Mobasher, B., Chartab, N., et al. 2023, ApJ, 951, 147
- Simard et al. (2002) Simard, L., Willmer, C. N. A., Vogt, N. P., et al. 2002, The Astrophysical Journal Supplement Series, 142, 1
- Skilling & Bryan (1984) Skilling, J. & Bryan, R. K. 1984, MNRAS, 211, 111
- Sok et al. (2022) Sok, V., Muzzin, A., Jablonka, P., et al. 2022, The Astrophysical Journal, 924, 7
- Starck et al. (2015) Starck, J.-L., Murtagh, F., & Bertero, M. 2015, Starlet Transform in Astronomical Data Processing, ed. O. Scherzer (New York, NY: Springer New York), 2053–2098
- Starck et al. (1995) Starck, J.-L., Murtagii, F., & Bijaoui, A. 1995, Graphical Models and Image Processing, 57, 420
- Sureau et al. (2020) Sureau, F., Lechat, A., & Starck, J.-L. 2020, A&A, 641, A67
- Tikhonov & Arsenin (1977) Tikhonov, A. N. & Arsenin, V. Y. 1977, Solutions of ill-posed problems (Washington, D.C.: John Wiley & Sons, New York: V. H. Winston & Sons), xiii+258, translated from the Russian, Preface by translation editor Fritz John, Scripta Series in Mathematics
- Vaswani et al. (2017) Vaswani, A., Shazeer, N., Parmar, N., et al. 2017, in Advances in Neural Information Processing Systems, ed. I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, & R. Garnett, Vol. 30 (Curran Associates, Inc.)
- Wang et al. (2004) Wang, Z., Bovik, A., Sheikh, H., & Simoncelli, E. 2004, IEEE Transactions on Image Processing, 13, 600
- Wang & Bovik (2009) Wang, Z. & Bovik, A. C. 2009, IEEE Signal Processing Magazine, 26, 98
- Wang et al. (2022) Wang, Z., Cun, X., Bao, J., et al. 2022, in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, 17683–17693
- White, S. D. M. et al. (2005) White, S. D. M., Clowe, D. I., Simard, L., et al. 2005, A&A, 444, 365
- Wuyts et al. (2012) Wuyts, S., Förster Schreiber, N. M., Genzel, R., et al. 2012, ApJ, 753, 114
- Yan et al. (2020) Yan, Q., Zhang, L., Liu, Y., et al. 2020, IEEE Transactions on Image Processing, 29, 4308
- Yu et al. (2019) Yu, S., Park, B., & Jeong, J. 2019, in Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition Workshops
- Yuan et al. (2021) Yuan, L., Chen, Y., Wang, T., et al. 2021, Tokens-to-Token ViT: Training Vision Transformers from Scratch on ImageNet
- Zamir et al. (2022) Zamir, S. W., Arora, A., Khan, S., et al. 2022, in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, 5728–5739
Appendix A Supplementary Figures
As a validity check, we ensured that any insights into the internal properties of our sample galaxies remain unaffected by size biases, whether concerning redshift or disc colors. This is illustrated in Figure 11, where (a) represents the histogram of galaxy sizes grouped by their parent cluster redshift, and (b) by disc colour. Both plots indicate that all galaxies share the same global spatial extent. Each individual legend in the histograms is normalized such that its probability sums up to one.
Appendix B Impact of debiasing with multi-resolution support on neural networks
For a more rigorous study of the impact of debiasing with multi-resolution support on neural networks, we considered three different neural networks with and without the multi-resolution debiasing: Learnlet (Ramzi et al. 2023), Unet-64 (Ronneberger et al. 2015), and SUNet (Fan et al. 2022). To provide a quantitative comparison, we computed the Structural Similarity Index Measure (SSIM) and the flux error in small-scale structures between galaxies extracted from CANDELS and their simulated degraded versions, as done in Akhaury et al. (2022). The selection of galaxies based on their FWHM and magnitude is depicted in Figure 12. To prevent the background noise in the HST images from biasing our metrics, we fit a Gaussian window around each object with an FWHM equal to its catalogue-derived value. The trends in the metrics, depicted in Figure 13, as a function of the object magnitude, indicate an enhancement in SSIM and flux error for all three networks after debiasing.
Appendix C Hallucinations and the impact of training loss function
To estimate the occurrence of unexpected artefacts or hallucinations introduced by neural networks, we applied the SCARLET detection procedure (as outlined in Section 4.1) with a tight detection threshold to each scale. To determine the number of false positives, we examined whether the centroid of a detection in the neural network’s output fell within a -pixel radius of the HST detection centroid. Detections failing to meet this criterion were considered false positives. Figure 14 illustrates the impact of different loss functions on the hallucination rate as a function of the galaxy FWHM and magnitude for three neural networks: Learnlet, Unet-64, and SUNet. We conducted the experiments on the same test dataset of 2232 galaxies as detailed in Appendix B. Notably, Unet-64 consistently exhibited improved performance when trained with the -loss.

### v2

@@
 As ground-based all-sky astronomical surveys will gather millions of images in the coming years, a critical requirement emerges for the development of fast deconvolution algorithms capable of efficiently improving the spatial resolution of these images.
-By successfully recovering clean and high-resolution images from these surveys, our objective is to help deepen our understanding of galaxy formation and evolution through accurate photometric measurements.
+By successfully recovering clean and high-resolution images from these surveys, the objective is to deepen the understanding of galaxy formation and evolution through accurate photometric measurements.
 Methods.
@@
 Results.
-By conducting a performance comparison between our deep learning-based method and Firedec, a classical deconvolution algorithm, we analyze a subset of the EDisCS cluster samples.
-We demonstrate the advantage of our method in terms of resolution recovery, generalization to different noise properties, and computational efficiency.
-Not only does the analysis of this cluster sample assess the efficiency of our method, but it also enables us to quantify the number of clumps within these galaxies in relation to their disc colour.
-This robust technique holds promise for identifying structures in the distant universe from ground-based images.
+We conducted a performance comparison between our deep learning-based method and Firedec, a classical deconvolution algorithm, based on an analysis of a subset of the EDisCS cluster samples.
+We demonstrate the advantage of our method in terms of resolution recovery, generalisation to different noise properties, and computational efficiency.
+The analysis of this cluster sample not only allowed us to assess the efficiency of our method, but it also enabled us to quantify the number of clumps within these galaxies in relation to their disc colour.
+This robust technique that we propose holds promise for identifying structures in the distant universe through ground-based images.
 Key Words.:
@@
 High spatial resolution and high signal-to-noise observations are prerequisites to most observational astrophysical problems.
-However, expecting the two conditions to happen simultaneously is challenging: space telescopes have limited collecting power, while large telescopes are ground-based and are therefore affected by atmospheric turbulence.
-This is clearly illustrated by the two main missions/observatories of the next decade: the ESA-NASA Euclid space telescope (Euclid Collaboration et al., 2022; Laureijs et al., 2011) and the Vera C.
+However, expecting the two conditions to happen simultaneously is challenging, as space telescopes have a limited collecting power, while large telescopes are ground-based and are therefore affected by atmospheric turbulence.
+This is clearly illustrated by the two main missions of the next decade: the ESA-NASA Euclid space telescope (Euclid Collaboration et al., 2022; Laureijs et al., 2011) and the Vera C.
 Rubin Observatory (Ivezić et al., 2019).
 Exploiting the best of both worlds is possible, provided reliable post-processing techniques are developed to remove blurring by the atmosphere and instrument point spread function (PSF).
-To further complicate matters, sensor variations introduce noise in the images.
-Hence, image deconvolution in astrophysics is an ill-posed and ill-conditioned inverse problem that requires regularization to achieve a unique solution.
-This was realized very early in the field, and solutions were proposed, such as minimizing the Tikhonov function (e.g.
+To further complicate matters, sensor variations introduce noise into images.
+Hence, image deconvolution in astrophysics is an ill-posed and ill-conditioned inverse problem that requires regularisation to achieve a unique solution.
+This was realised very early in the field, and solutions were proposed, such as minimising the Tikhonov function (e.g.
 Tikhonov & Arsenin, 1977) or maximizing the entropy of the solution (e.g.
 Skilling & Bryan, 1984).
-Other algorithms, based on Bayesian statistics, include the Lucy-Richardson algorithm, used on the early data from the Hubble Space Telescope (HST) (Richardson, 1972; Lucy, 1974).
+Other algorithms, based on Bayesian statistics, include the Lucy-Richardson algorithm, used on the early data from the Hubble Space Telescope (HST; Richardson (1972); Lucy (1974)).
 Magain et al.
-(MCS; 1998) proposed a two-channel method that separates the point sources from the spatially extended ones and that deconvolves using a narrow PSF, hence achieving finite but improved resolution well compatible with the sampling (pixel) chosen to represent the solution.
-Improvements of the ”MCS” method implement wavelet regularization of the extended channel of the solution Cantale et al.
-(2016a).
-This method was further refined by Michalewicz et al.
-(STARRED; 2023) who use an isotropic wavelet basis, called Starlets (Starck et al., 2015), to regularize the solution.
+(MCS; 1998) proposed a two-channel method that separates the point sources from the spatially extended ones and deconvolves using a narrow PSF, hence achieving finite but improved resolution compatible with the sampling (pixel) chosen to represent the solution.
+Improvements of the ‘MCS’ method implement wavelet regularisation of the extended channel of the solution Cantale et al.
+(2016a), and this improved method was further refined by Michalewicz et al.
+(STARRED; 2023), who used an isotropic wavelet basis called Starlets (Starck et al., 2015) to regularise the solution.
 However, deep learning offers a completely different approach, by learning the properties of the desired solution.
@@
 (2020) developed the Tikhonet method for deconvolving galaxy images in the optical domain.
-They demonstrated that Tikhonet outperformed sparse regularization-based methods in terms of Mean Squared Error (MSE) and a shape criterion, where a measure of the galaxy ellipticity was used to encode its shape.
+They demonstrated that Tikhonet outperformed sparse regularisation-based methods in terms of mean squared error and a shape criterion, where a measure of the galaxy ellipticity was used to encode its shape.
 Nammour et al.
 (2022) added a shape constraint to the Tikhonet loss function and achieved better performance.
-In our recent work (Akhaury et al., 2022), we proposed a new deconvolution approach that employs the Learnlet decomposition (Ramzi et al., 2023), using the same two-step approach as in Sureau et al.
-(2020), but by substituting the Unet denoiser by Learnlet.
+In our recent work (Akhaury et al., 2022), we have proposed a new deconvolution approach that employs the Learnlet decomposition (Ramzi et al., 2023).
+It uses the same two-step approach as in Sureau et al.
+(2020) but substitutes the UNet denoiser by Learnlet.
 With the recent advent of Vision Transformers (Dosovitskiy et al., 2021), significant progress has been made in the field of image restoration (Liang et al., 2021; Zamir et al., 2022; Wang et al., 2022).
@@
 The bias appears as a small flux loss in the outputs.
-To counter this, we propose a third debiasing procedure based on the active coefficients in the sparsity wavelet framework, called Multi-resolution support (Starck et al., 1995), as explained in Section 2.3.
-Our experiments involved real HST images, with the network trained on images extracted from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011).
-Finally, to assess the generalizability, we test our new code on a completely different sample of ground-based images obtained with the FORS2 camera at the Very Large Telescope (VLT).
+To counter this, we propose a third debiasing procedure based on the active coefficients in the sparsity wavelet framework called multi-resolution support (Starck et al., 1995), explained in Section 2.3.
+Our experiments involved real HST images, and the network was trained on images extracted from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011).
+Finally, to assess its generalisability, we tested our new code on a completely different sample of ground-based images obtained with the FORS2 camera at the Very Large Telescope (VLT).
 Spatially resolved images of galaxies can help address a multitude of topics involving their morphologies.
-In the present case, we tackle the question of how the galaxy cluster environment could impact the properties of the discs of spiral galaxies.
+In the present case, we tackle the question of how the galaxy cluster environment can impact the properties of the discs of spiral galaxies.
 From the same dataset, Cantale et al.
-(2016b) had found that of cluster spiral galaxies at redshift - had disc colours redder by more than of the mean colours of their field counterparts.
+(2016b) found that of cluster spiral galaxies at redshift - have disc colours that are redder by more than of the mean colours of their field counterparts.
 This result was obtained thanks to the deconvolution code Firedec (Cantale et al., 2016a).
-The VLT and -band images, with initial spatial resolutions between and had been deconvolved with a target final resolution of on pixels.
+The VLT - and -band images, with initial spatial resolutions between and , were deconvolved with a target final resolution of on pixels.
 Even though the gain in resolution was indeed substantial, conditions were not sufficient to go beyond the global photometric properties of the discs.
 In this study, we go one step further and investigate their internal structure, specifically by identifying their star-forming clumps.
-Compact star-forming clumps are identified in distant galaxies, particularly when they benefit from HST deep images (e.g., Wuyts et al., 2012; Guo et al., 2015; Sattari et al., 2023).
-They are understood as playing a crucial role in galaxy assembly and star-formation activity.
+Compact star-forming clumps have been identified in distant galaxies, particularly with the aid of HST deep images (e.g., Wuyts et al., 2012; Guo et al., 2015; Sattari et al., 2023).
+They are understood to play a crucial role in galaxy assembly and star-formation activity.
 Recent research by Sok et al.
-(2022) explored clump fractions in star-forming galaxies (SFGs) from multi-band analysis in the COSMOS field.
+(2022) explored clump fractions in star-forming galaxies from multi-band analysis in the COSMOS field.
 The match between HST and ground-based resolution was performed with Firedec.
@@
 Finally, we draw our conclusions in Section 5.
-To support reproducible research, the codes and trained models utilized in this article are publicly accessible in Section 6.
+To support reproducible research, the codes and trained models utilised in this article are publicly accessible in Section 6.
 Additional studies and supplementary information can be found in Appendices A, B, and C.
-2 Deep-Learning based Deconvolution
-The deconvolution problem can be summarized in a very simple (but hard to solve) equation. Let be the observed image and be the PSF. The observed image can be modelled as
+2 Deep learning-based deconvolution
+The deconvolution problem can be summarised with a very simple (but hard to solve) equation. Let be the observed image and be the PSF. The observed image can be modelled as
 where denotes the target image, denotes the convolution operation, and denotes additive Gaussian noise.
-The goal is then to recover the ground truth image , given the PSF convolution and the unknown noise.
-Such ill-posed inverse problems require regularization of the solution, to select the most appropriate one among the many others that are compatible with the data.
-Sparse wavelet regularization using the or norm has remained the most accepted one in the past, but the recent advent of machine learning methods has changed this paradigm.
+The goal is to recover the ground truth image , given the PSF convolution and the unknown noise.
+Such ill-posed inverse problems require regularisation of the solution in order to select the one that is most appropriate compared to the many others that are compatible with the data.
+Sparse wavelet regularisation using the or norm remained the most accepted regularisation in the past, but the recent advent of machine learning methods has changed the paradigm.
 2.1 Tikhonov deconvolution
 Tikhonov Deconvolution is a two-step deep learning-based deconvolution technique.
-In the first step, the input images undergo deconvolution using a Tikhonov filter with quadratic regularization.
-If denotes the circulant matrix associated with the convolution operator , the Tikhonov solution of equation 1 is expressed as:
-Here, represents the linear Tikhonov filter, configured as a Laplacian high-pass filter to penalize high frequencies.
-The regularization weight, denoted as , is determined through a grid search.
+In the first step, the input images undergo deconvolution using a Tikhonov filter with quadratic regularisation.
+If denotes the circulant matrix associated with the convolution operator , the Tikhonov solution of equation 1 is expressed as
+Here, represents the linear Tikhonov filter, configured as a Laplacian high-pass filter to penalise high frequencies.
+The regularisation weight, denoted as , is determined through a grid search.
 The Tikhonov step produces deconvolved images containing correlated additive noise, which is subsequently eliminated in the following step by an appropriate denoiser.
-The denoisers are trained to learn the mapping from the Tikhonov output to the ground truth image by minimizing a suitable loss function, such as or .
+The denoisers are trained to learn the mapping from the Tikhonov output to the ground truth image by minimising a suitable loss function, such as or .
 The denoising performance is significantly influenced by the choice of the model architecture.
@@
 Additionally, Mohan et al.
-(2020) demonstrated that biases in convolutional layers can result in low generalization capability.
-Consequently, for our experiments, we opt for bias-free networks.
+(2020) demonstrated that biases in convolutional layers can result in a low generalisation capability.
+Consequently, for our experiments, we opted for bias-free networks.
 2.2 SUNet denoising
-In recent times, the UNet architecture has become popular in various image-processing applications due to its incorporation of hierarchical feature maps, facilitating the acquisition of multi-scale contextual features.
+In recent times, the UNet architecture has become popular in various image-processing applications due to its incorporation of hierarchical feature maps, which facilitate the acquisition of multi-scale contextual features.
 It is widely employed in diverse computer vision tasks, including segmentation and restoration (Yu et al., 2019; Gurrola-Ramos et al., 2021).
 Evolved versions such as Dense-UNet (Guan et al., 2020), Res-UNet (nan Xiao et al., 2018), Non-local UNet (Yan et al., 2020), and Attention UNet (Jin et al., 2020) have also emerged.
-Thanks to its flexible structure, the UNet can adapt to different building blocks, enhancing its overall performance.
-Moreover, the evolution of image processing methodologies has seen the introduction of Transformer models (Vaswani et al., 2017), which were initially successful in natural language processing (NLP) but have also demonstrated impressive performance in image classification (Dosovitskiy et al., 2021; Yuan et al., 2021).
-However, when directly applied to vision tasks, transformers face challenges such as the large-scale difference between images and sequences, making them less effective in modelling long sequences due to the need for a square number of parameters for 1-dimensional sequences.
-Additionally, transformers are not well-suited for dense prediction tasks such as instance segmentation at the pixel level.
+Thanks to its flexible structure, UNet can adapt to different building blocks, enhancing its overall performance.
+Moreover, the evolution of image-processing methodologies has seen the introduction of transformer models (Vaswani et al., 2017), which were initially successful in natural language processing but have also demonstrated impressive performance in image classification (Dosovitskiy et al., 2021; Yuan et al., 2021).
+However, when directly applied to vision tasks, transformers face challenges such as the large-scale difference between images and sequences, making them less effective in modelling long sequences due to a need for a square number of parameters for one-dimensional sequences.
+Additionally, transformers are not well suited for dense prediction tasks such as instance segmentation at the pixel level.
 Swin Transformer addresses these issues by introducing a shifted-window mechanism to reduce the number of parameters, establishing itself as a state-of-the-art solution for high-level vision tasks (Liu et al., 2021).
@@
 The network architecture is visually depicted in Figure 1, and the PyTorch code is publicly available on GitHub (details in Section 6).
-While SUNet was originally developed for White Gaussian Noise (WGN) removal, the Tikhonov deconvolution step (equation 2) introduces alterations to the noise characteristics in the images.
-Consequently, it becomes crucial to assess the extent to which SUNet can effectively handle the presence of Colored Gaussian Noise (CGN).
-Thus, our tests examine SUNet’s generalizability in the presence of CGN.
+While SUNet was originally developed for white Gaussian noise removal, the Tikhonov deconvolution step (equation 2) introduces alterations to the image’s noise characteristics.
+Consequently, it becomes crucial to assess the extent to which SUNet can effectively handle the presence of coloured Gaussian noise.
+Thus, our tests examine SUNet’s generalisability in the presence of CGN.
 2.3 Debiasing with multi-resolution support
 Our study reveals that the deep learning-based solution introduces a bias, evident in the form of positive structures in the residuals, as illustrated in Figure 2(a).
-This bias can have implications for the accuracy of scientific analyses, potentially influencing the flux estimation within the recovered features in the reconstructed images.
-To recover the lost flux and enhance image sharpness, multi-resolution support has proven effective (Starck et al., 1995).
-Denoting the Starlet transform as and the SUNet output solution as , an iterative process allows the recovery of flux from the residual .
-In each iteration, denoted as , a debiasing correction term is applied by multiplying the multi-resolution support matrices of the SUNet output solution at each scale with the Starlet decomposition of the gradient of the residual, incrementally modifying the deconvolved image:
+This bias can have implications on the accuracy of the scientific analyses, potentially influencing the flux estimation within the recovered features in the reconstructed images.
+In recovering the lost flux and enhancing image sharpness, multi-resolution support has been proven effective (Starck et al., 1995).
+Denoting the Starlet transform as and the SUNet output solution as , an iterative process allowed for the recovery of flux from the residual .
+In each iteration, denoted as , a debiasing correction term was applied by multiplying the multi-resolution support matrices of the SUNet output solution at each scale with the Starlet decomposition of the gradient of the residual, incrementally modifying the deconvolved image:
 where
-The acronym MRS stands for multi-resolution support, indicating a boolean measure of whether an image contains information at a specific pixel and scale.
+, and
+The acronym ‘MRS’ stands for multi-resolution support and indicates a Boolean measure of whether an image contains information at a specific pixel and scale.
 If represents the wavelet coefficient at a given scale and is the threshold value, the multi-resolution support operation can be defined as:
 The flux-recovery process is illustrated in Figure 2.
-The iterative process is stopped once convergence is achieved in the standard deviation of the residual as a function of the number of iterations, as seen in sub-figure 2(d).
+The iterative process was stopped once convergence was achieved in the standard deviation of the residual as a function of the number of iterations, as seen in sub-figure 2(d).
 A more detailed study on the impact of debiasing with multi-resolution support on neural networks is presented in Appendix B.
@@
 3.1 Training dataset generation
-We extracted Hubble Space Telescope (HST) cutouts measuring pixels from CANDELS (Grogin et al., 2011; Koekemoer et al., 2011) in the filter (-band).
+We extracted HST cutouts measuring pixels from CANDELS (Grogin et al., 2011; Koekemoer et al., 2011) in the filter (-band).
 These cutouts were then convolved with a Gaussian PSF having a full width at half maximum (FWHM) of 15 pixels.
-Following the convolution, we injected White Gaussian Noise (WGN) with a standard deviation denoted as .
-This choice ensured that the faintest object in our dataset had a peak Signal-to-Noise Ratio (SNR) close to and was barely visible.
-With this particular value of , our dataset exhibited a range of SNR values depending on the magnitude of each galaxy.
-To standardize the images, each one was normalized within the range.
-This normalization involved subtracting its own mean and scaling its peak value by the maximum value in the image.
+Following the convolution, we injected white Gaussian noise with a standard deviation denoted as .
+This choice ensured that the faintest object in our dataset had a peak signal-to-noise (S/N) close to one and was barely visible.
+With this particular value of , our dataset exhibited a range of S/N values depending on the magnitude of each galaxy.
+To standardise the images, each image was normalised within the range.
+This normalization involved subtracting the image’s mean and scaling the peak value by the image’s maximum value.
 Finally, the batch of images was randomly divided into training-validation-test subsets in the ratio .
@@
 The training aimed at learning the mapping from the Tikhonov output (containing CGN) to the corresponding HST image .
-The training utilized the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of , gradually halving every 25 epochs until reaching a minimum of .
-As in the original SUNet paper (Fan et al., 2022), the -loss was used for training.
+The training utilised the Adam optimiser (Kingma & Ba, 2014) with an initial learning rate of , gradually halving every 25 epochs until reaching a minimum of .
+As in the original SUNet paper (Fan et al., 2022), -loss was used for training.
 A more detailed discussion on the significance of the training loss function is given in Appendix C.
 The input images were processed in mini-batches of size .
-The dataset was augmented with random rotations in multiples of 90°, translations and flips along horizontal & vertical axes.
-Starting with images in our initial training dataset, we increase its diversity by a factor of by performing augmentation.
+The dataset was augmented with random rotations in multiples of 90°; translations and flips were along horizontal and vertical axes.
+Starting with images in our initial training dataset, we increased its diversity by a factor of ten by performing augmentation.
 3.3 Test dataset
-We utilized the ESO Distant Cluster Survey, EDisCS (White, S.
+We utilised the ESO Distant Cluster Survey (EDisCS; White, S.
 D.
 M.
-et al., 2005), as our benchmark dataset to assess the performance of our deconvolution method.
-EDisCS is an extensive ESO Large Programme focused on the analysis of galaxy clusters within the redshift range , covering a diverse range of masses with velocity dispersions ranging from approximately km s-1 to km s-1.
-All clusters benefit from deep , , , and photometry obtained with FORS2 at the VLT.
+et al., 2005) as our benchmark dataset to assess the performance of our deconvolution method.
+EDisCS is an extensive ESO Large Programme focused on the analysis of galaxy clusters within the redshift range and covering a diverse range of masses, with velocity dispersions ranging from approximately km s-1 to km s-1.
+All of the clusters benefited from the deep , , , and photometry obtained with FORS2 at the VLT.
 Additionally, a subset of ten clusters was imaged with the ACS at the HST in the F814W filter (Simard et al., 2002).
 Previous work by Cantale et al.
-(2016b) employed the deconvolution technique Firedec to analyze spiral disc colours for EDisCS cluster galaxies, investigating trends with cluster masses and lookback-time.
-In our study, we focus on analyzing a subset of EDisCS clusters at three distinct redshifts—, , and —using our proposed deconvolution method, and go a step further by investigating the star-forming clumps in these galaxies.
+(2016b) employed the deconvolution technique Firedec to analyse spiral disc colours for EDisCS cluster galaxies, investigating trends with cluster masses and lookback time.
+For our study, we focused on analysing a subset of EDisCS clusters at three distinct redshifts—, , and —using our proposed deconvolution method, and we went a step further by investigating the star-forming clumps in these galaxies.
 Table 1 provides a summary of the properties of these clusters and the number of galaxies in each of them.
@@
 4 Results
-The EDisCS images in the , , and bands served as inputs for the SUNet deconvolution framework, yielding corresponding deconvolved outputs.
-Remarkably, the algorithm operates with exceptional speed, requiring only ms to deconvolve a single image on a Titan RTX Turing GPU with 24 GB RAM—approximately times faster than Firedec on average.
+The EDisCS images in the -, -, and -bands served as inputs for the SUNet deconvolution framework, yielding corresponding deconvolved outputs.
+The algorithm operates with exceptional speed, requiring only ms to deconvolve a single image on a Titan RTX Turing GPU with 24 GB RAM—approximately times faster than Firedec, on average.
 A histogram of the magnitudes of the same objects in the HST F814W filter is shown in Figure 3.
@@
 Section 4.2 then presents a comparative analysis of the quality of our SUNet outputs against Firedec (Cantale et al., 2016a).
-With the validity of our method established, Section 4.3 presents the deconvolved results, followed by a thorough analysis in Section 4.4.
+With the validity of our method established, Section 4.3 presents the deconvolved results and is followed by a thorough analysis in Section 4.4.
 4.1 Objects size and clump detection
 To detect clumps, or small-scale structures, and measure the sizes of galaxies, we employed the SCARLET Python package (Melchior et al., 2018).
-Utilizing the Starlet transform from SCARLET, our approach involved decomposing images into different scales, where each scale captures a specific frequency component.
-Ensuring consistency for an unbiased comparison, we maintained fixed algorithm parameters across different bands and objects.
+Utilising the Starlet transform from SCARLET, our approach involved decomposing images into different scales, where each scale captures a specific frequency component.
+To ensure consistency for an unbiased comparison, we maintained fixed algorithm parameters across different bands and objects.
 All images underwent decomposition into five scales, with the fourth scale chosen for size detection and the second scale for clump detection.
@@
 4.2 Comparison with classical methods
-We conducted a thorough performance comparison between SUNet and Firedec, a classical deconvolution method based on wavelet regularization (Cantale et al., 2016a).
-For direct comparison with HST quality, we concentrate on the -band outputs for each method, as the EDisCS clusters were exclusively observed in the F814W filter for HST.
-Both methods exhibit better performance on low-magnitude (or high-SNR) images, with a gradual decline in performance in the high-magnitude (or low-SNR) regime.
-In this case, the Mean Squared Error (MSE) metric between the deconvolved outputs and the ground truth HST images is not a robust metric for indicating similarity, since it would be biased by the background noise in the HST images (Wang & Bovik, 2009).
-Instead, we use the Structural Similarity Index (SSIM), a full reference metric which quantifies the similarity between two images by comparing their structural information or spatial inter-dependencies (Wang et al., 2004).
-An SSIM of 1 implies identical images.
+We conducted a thorough performance comparison between SUNet and Firedec, a classical deconvolution method based on wavelet regularisation (Cantale et al., 2016a).
+For direct comparison with HST quality, we concentrated on the -band outputs for each method, as the EDisCS clusters were exclusively observed in the F814W filter for HST.
+Both methods exhibited a better performance on low-magnitude (or high-S/N) images, with a gradual decline in performance in the high-magnitude (or low-S/N) regime.
+In this case, the mean squared error metric between the deconvolved outputs and the ground truth HST images is not a robust metric for indicating similarity since it is biased by the background noise in the HST images (Wang & Bovik, 2009).
+Instead, we used the structural similarity index measure (SSIM), a full reference metric which quantifies the similarity between two images by comparing their structural information or spatial interdependencies (Wang et al., 2004).
+An SSIM of one implies identical images.
 The observed trends are illustrated in Figure 5.
@@
 (2023).
-STARRED brings innovation by incorporating an isotropic wavelet basis known as Starlets (Starck et al., 2015) to refine the regularization process in solving the deconvolution problem.
+STARRED brings innovation by incorporating an isotropic wavelet basis known as Starlets (Starck et al., 2015) that can refine the regularisation process when solving the deconvolution problem.
 The outputs of Firedec, STARRED, and SUNet are shown in Figure 7.
-Upon visual inspection of the outputs and the residuals (residual noisy image PSF deconvolved image), it is evident that SUNet consistently generalizes better than Firedec and STARRED.
+Upon visual inspection of the outputs and the residuals (residual noisy image PSF deconvolved image), it was evident that SUNet consistently generalises better than Firedec and STARRED.
 4.3 SUNet deconvolution results
-The methods were put to the test using real ground-based images captured by the FORS2 camera at the Very Large Telescope (VLT) in Chile.
-Notably, SUNet showed a remarkable ability to generalize effectively to images with entirely different noise properties than those present in the training dataset, as depicted in Figure 8.
-As illustrated, we are successfully able to recover the morphology and lost small-scale structures.
+The methods were put to the test using real ground-based images captured by the FORS2 camera at VLT in Chile.
+Notably, SUNet showed a remarkable ability to effectively generalise to images with entirely different noise properties than those present in the training dataset, as depicted in Figure 8.
+As illustrated, we were able to successfully recover the morphology and lost small-scale structures.
 Figure 9 presents the trend in the relative flux error between the SUNet deconvolved outputs and the corresponding HST targets as a function of the total clump size.
@@
 To gauge the achieved resolution in the deconvolved outputs, we calculated the average ratio between the areas of the smallest detected clump in the SUNet output and its counterpart in the HST image.
-This ratio, approximately , implies an average SUNet output resolution of around , considering the known HST resolution of .
+This ratio, approximately , implies an average SUNet output resolution of around , considering the known HST resolution is .
 4.3.2 False positives and false negatives
@@
 This rate was found to be , signifying a very low probability of missing features.
+To obtain a more statistically robust evaluation, we tested the method on another dataset of galaxies extracted from CANDELS (Grogin et al., 2011) and compared it with other neural networks.
+This work is shown in Appendix C.
 4.4 Analysis of deconvolved EDisCS galaxies
 As a sanity check, we verified that any conclusions drawn regarding the internal properties of our sample galaxies are not influenced by biases in their sizes, either in relation to redshift or disc colours.
-To this end, the histogram of galaxy sizes grouped by their parent cluster redshift and disc colour is presented in Appendix A.
+To this end, a histogram of galaxy sizes grouped by their parent cluster redshift and disc colour is presented in Appendix A.
 Both plots show that all galaxies have the same global spatial extent.
@@
 (2016b), our analysis focuses on the population of galaxies with discs redder than their field counterparts.
-Some of our sample galaxies have normal colours, hence falling in the colour distribution of the field galaxies, some are bluer (by more than of the colour distribution).
-A few known physical processes can induce enhanced star-forming activity, but we are instead interested in the possible evidence for quenching mechanisms.
-Therefore, in the latter case, normal and blue disc galaxies form a common broad class of systems to which we compare the redder ones.
-Employing the clump detection method outlined in Section 4.1, we computed the histogram of the number of clumps in galaxies, categorized by disc colour, as shown in Figure 10.
-Note that the only-1-clump case reflects the identification of the central and luminous part (bulge) of the galaxies.
-As illustrated in Figures 4 and 8, clumps in the -band are brighter than those in the and -bands, in agreement with the spectral energy distribution of young stellar populations.
+Some of our sample galaxies have normal colours and hence fall into the colour distribution of the field galaxies, and some are bluer (by more than of the colour distribution).
+A few known physical processes can induce enhanced star-forming activity, but we were instead interested in the possible evidence for quenching mechanisms.
+Therefore, in the latter case, normal and blue disc galaxies form a common broad class of systems to which we compared the redder ones.
+Employing the clump detection method outlined in Section 4.1, we computed the histogram of the number of clumps in galaxies, categorised by disc colour, as shown in Figure 10.
+We note that the only-one-clump case reflects the identification of the central and luminous part (bulge) of the galaxies.
+As illustrated in Figures 4 and 8, clumps in the -band are brighter than those in the - and -bands, in agreement with the spectral energy distribution of young stellar populations.
 In principle, it is therefore easier to detect clumps in the -band at equivalent photometric depths of the images.
-This may explain the more continuous distribution in the number of clumps, from to , in the -band.
-Even so, as witnessed in Figure 10, the general trend is the same from one band to the other.
-This trend is clear and reveals that the red discs, that were initially identified by Cantale et al.
-(2016b), have fewer clumps than their bluer counterparts, most likely due to an earlier cessation of star formation.
-This result opens promising prospects for future studies on larger samples and over larger look-intervals.
+This may explain the more continuous distribution in the number of clumps, from one to six, in the -band.
+Even so, as we witnessed in Figure 10, the general trend is the same from one band to the other.
+This trend is clear and reveals that the red discs that were initially identified by Cantale et al.
+(2016b) have fewer clumps than their bluer counterparts, most likely due to an earlier cessation of star formation.
+This result opens promising prospects for future studies on larger samples and over larger look intervals.
 5 Conclusion
-We proposed a deconvolution framework involving a two-step process: Tikhonov deconvolution, post-processing with an SUNet denoiser, and an additional debiasing step using multi-resolution support.
+We have proposed a deconvolution framework involving a two-step process—namely, Tikhonov deconvolution and post-processing with an SUNet denoiser—and an additional debiasing step using multi-resolution support.
 SUNet was trained on galaxy images from the CANDELS survey and demonstrated superior performance compared to Firedec in the astrophysical context.
@@
 We quantified the number of clumps in these galaxies, examining their relationship with disc colour.
-Our results, based on both quantitative metrics and visual assessments, highlighted the effectiveness of SUNet, showcasing its ability to generalize well to unseen real images with diverse noise properties, which can be attributed to its transformer-based backbone involving the self-attention mechanism.
-In summary, this work introduces and evaluates an advanced deconvolution framework applied to ground-based astronomical images. The key findings and contributions include:
-Resolution recovery: Based on our SCARLET detection procedure, SUNet demonstrates the capability to recover small-scale structures with an average resolution of approximately and outperforms classical algorithms like STARRED and Firedec (Section 4.3.1).
-Generalization to diverse noise properties: The method showcases robust generalization to noise properties different from its training dataset, indicating its adaptability to various observational conditions.
+Our results, based on both quantitative metrics and visual assessments, highlight the effectiveness of SUNet and showcase its ability to generalise unseen real images with diverse noise properties, which can be attributed to its transformer-based backbone involving the self-attention mechanism.
+In summary, this work introduces and evaluates an advanced deconvolution framework applied to ground-based astronomical images. The key findings and contributions include the following:
+Resolution recovery: Based on our SCARLET detection procedure, SUNet demonstrates the capability to recover small-scale structures, with an average resolution of approximately , and it outperforms classical algorithms such as STARRED and Firedec (Section 4.3.1).
+Generalisation to diverse noise properties: The method showcases robust generalisation to noise properties different from its training dataset, indicating its adaptability to various observational conditions.
 Clump analysis: Red discs exhibit fewer clumps than their bluer counterparts, affirming the lower presence of star-forming regions.
 False positive and false negative rates: Based on our SCARLET detection analysis on EDisCS, SUNet maintains a false positive rate of and a false negative rate of , ensuring reliable feature recovery (Section 4.3.2).
-Computational efficiency: The AI-based framework proves to be highly efficient, with an execution time of approximately ms per image, making it around times faster than traditional deconvolution methods like STARRED and Firedec.
-Hence, this technique can be used to efficiently identify structures in the distant universe at high spatial resolution from ground-based images.
+Computational efficiency: The AI-based framework proves to be highly efficient, with an execution time of approximately ms per image, making it around times faster than traditional deconvolution methods such as STARRED and Firedec.
+Our proposed technique can therefore be used with ground-based images to efficiently identify structures in the distant universe at high spatial resolution.
 The technique’s applicability to multi-band observations further enhances its utility in studying various astrophysical phenomena.
@@
 Access to such a fast and robust deconvolution framework holds the potential to facilitate numerous astrophysical investigations.
-6 Reproducible Research
-[…]
+6 Data availability
+For the sake of reproducible research, the codes and the trained models used for this article are publicly available online.
 1.
-[…]
+The ready-to-use version of our deconvolution method222https://github.com/utsav-akhaury/SUNet/tree/main/Deconvolution.
 2.
-[…]
+The repository fork of the SUNet code used for training the network333https://github.com/utsav-akhaury/SUNet.
 3.
-[…]
+Link to the trained network weights444https://doi.org/10.5281/zenodo.10287213.
 4.
-[…]
+The repository forks of the Learnlet and Unet-64 codes used for comparison in the Appendices B and C555https://github.com/utsav-akhaury/understanding-unets/tree/candels.
 Acknowledgements.
@@
 - Zamir et al. (2022) Zamir, S. W., Arora, A., Khan, S., et al. 2022, in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, 5728–5739
-Appendix A Supplementary Figures
-As a validity check, we ensured that any insights into the internal properties of our sample galaxies remain unaffected by size biases, whether concerning redshift or disc colors.
-This is illustrated in Figure 11, where (a) represents the histogram of galaxy sizes grouped by their parent cluster redshift, and (b) by disc colour.
+Appendix A Supplementary figure
+As a validity check, we ensured that any insights into the internal properties of our sample galaxies remain unaffected by size biases, whether concerning redshift or disc colours.
+This is illustrated in Figure 11, where (a) represents the histogram of galaxy sizes grouped by their parent cluster redshift and (b) represents the same but grouped by disc colour.
 Both plots indicate that all galaxies share the same global spatial extent.
-Each individual legend in the histograms is normalized such that its probability sums up to one.
+Each individual legend in the histograms has been normalised such that its probability adds up to one.
 Appendix B Impact of debiasing with multi-resolution support on neural networks
@@
 2022).
-To provide a quantitative comparison, we computed the Structural Similarity Index Measure (SSIM) and the flux error in small-scale structures between galaxies extracted from CANDELS and their simulated degraded versions, as done in Akhaury et al.
+To provide a quantitative comparison, we computed the SSIM and the flux error in small-scale structures between galaxies extracted from CANDELS and their simulated degraded versions, as done in Akhaury et al.
 (2022).
@@
 To prevent the background noise in the HST images from biasing our metrics, we fit a Gaussian window around each object with an FWHM equal to its catalogue-derived value.
-The trends in the metrics, depicted in Figure 13, as a function of the object magnitude, indicate an enhancement in SSIM and flux error for all three networks after debiasing.
+The trends in the metrics, depicted in Figure 13 as a function of the object magnitude, indicate an enhancement in SSIM and flux error for all three networks after debiasing.
 Appendix C Hallucinations and the impact of training loss function
 To estimate the occurrence of unexpected artefacts or hallucinations introduced by neural networks, we applied the SCARLET detection procedure (as outlined in Section 4.1) with a tight detection threshold to each scale.
-To determine the number of false positives, we examined whether the centroid of a detection in the neural network’s output fell within a -pixel radius of the HST detection centroid.
+To determine the number of false positives, we examined whether the centroid of a detection in the neural network’s output fell within a five-pixel radius of the HST detection centroid.
 Detections failing to meet this criterion were considered false positives.
 Figure 14 illustrates the impact of different loss functions on the hallucination rate as a function of the galaxy FWHM and magnitude for three neural networks: Learnlet, Unet-64, and SUNet.
-We conducted the experiments on the same test dataset of 2232 galaxies as detailed in Appendix B.
+We conducted the experiments on the same test dataset of 2232 galaxies detailed in Appendix B.
 Notably, Unet-64 consistently exhibited improved performance when trained with the -loss.